OpenStudy (anonymous):

what are the excluded values of a/b-2

OpenStudy (anonymous):

\[\frac{ a }{ b } - 2 = \frac{ a }{ b } - \frac{ 2 }{ 1 } = \frac{ a - 2b }{ b }\]

OpenStudy (anonymous):

excluded values?